OpenStudy (anonymous):

Help! I'll give a medal!! Input Output 1 3 2 -1 3 -5 4 -9 5 ? ...n ? I have to complete the table...I can't figure out the pattern though, any suggestions?

OpenStudy (anonymous):

+2-2

OpenStudy (cwrw238):

look at the outputs they are an arithmetical sequence with common difference -4

OpenStudy (anonymous):

i think its -4 idk

OpenStudy (anonymous):

Ok the first blank would be -13 @cwrw238 ? right

OpenStudy (cwrw238):

the terms in an AP are a, a+d, a+ 2d + .. = a+(n-1)d

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

and nth term is a+ (n-1)d here a is the first term = 3 and d = -4

OpenStudy (anonymous):

ok, but I have to find what equation or pattern they use to get from the input to the output

OpenStudy (cwrw238):

the nth term gives the pattern plug in a=3 and d = -4

OpenStudy (anonymous):

a + (n - 1)d 3 + (n -1)-4 Is that how it should look? But I don't need to find n....

OpenStudy (cwrw238):

right - no just the nth term you can simplify that if you like nth term = 3 + (n-1)*-4 = 3 - 4n + 4 = 7 - 4n

OpenStudy (cwrw238):

we can check if this is correct try 5th term; = 7 - 4(5) = 7 - 20 = -13

OpenStudy (cwrw238):

6th term = 7 - 4(6) = -17 which is correct

OpenStudy (anonymous):

ohhhh I see....ok so the formula I use is 7 - 4n....so then the blank next to the nth input would b the formula 7-4n becuz I use that to find any output...right?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

awesome! thank you very much!!

OpenStudy (cwrw238):

yw